# Datasets for Fairness in Language Models: An In-Depth Survey

### v1

@@ Abstract
Fairness benchmarks play a central role in shaping how we evaluate language models, yet surprisingly little attention has been given to examining the datasets that these benchmarks rely on. This survey addresses that gap by presenting a broad and careful review of the most widely used fairness datasets in current language model research, characterizing them along several key dimensions including their origin, scope, content, and intended use to help researchers better appreciate the assumptions and limitations embedded in these resources. To support more meaningful comparisons and analyses, we introduce a unified evaluation framework that reveals consistent patterns of demographic disparities across datasets and scoring methods. Applying this framework to twenty four common benchmarks, we highlight the often overlooked biases that can influence conclusions about model fairness and offer practical guidance for selecting, combining, and interpreting these datasets. We also point to opportunities for creating new fairness benchmarks that reflect more diverse social contexts and encourage more thoughtful use of these tools going forward. All code, data, and detailed results are publicly available at https://github.com/vanbanTruong/Fairness-in-Large-Language-Models/tree/main/datasets to promote transparency and reproducibility across the research community.
@@ 1 Introduction
Language Models (LMs), such as BERT [1], ELMo [2], RoBERTa [3], GPT-4 [4], LLaMA-2 [5], LLaMa-3 [6], and BLOOM [7], have demonstrated remarkable performance in various natural language processing tasks, including translation [8], text sentiment analysis [9], and text summarization [10]. This success has led to their widespread adoption across high-stakes domains such as hiring [11], loan approvals [12], legal sentencing [13], and medical diagnoses [14]. However, as their influence grows, a critical concern has emerged: most LMs were not designed with fairness considerations, which could potentially lead to discriminatory outcomes in these downstream applications. For example, research has shown that these models can perpetuate stereotypes, linking leadership qualities with men while describing women as “warm and friendly” [15]. Such biases in LMs raise significant ethical and societal concerns, underscoring the need for rigorous fairness research in large language models.
@@
To address these fairness concerns, an expanding body of research focuses on understanding, assessing, and mitigating the biased behaviors exhibited by large language models. Datasets are central to these efforts, serving not only as benchmarks for identifying fairness issues but also as critical foundations for designing effective mitigation strategies [23]. For instance, WinoBias by Zhao et al. [55], which uses controlled sentence structures to isolate gender bias; CrowS-Pairs by Nangia et al. [56], which employs crowd-sourcing to capture naturally occurring stereotypes; and BBQ by Parrish et al. [57], which relies on expert annotation for high-quality bias measurements across nine demographic categories. These datasets provide controlled environments to systematically assess how models respond to different demographic groups, employing methods ranging from counterfactual approaches to specially designed prompts.
Despite the proliferation of these datasets, researchers face two critical challenges when selecting appropriate datasets from the numerous available options [23]. Firstly, determining which datasets best align with specific fairness goals is difficult due to the variability in dataset characteristics, inherent biases, and evaluation contexts. Secondly, bias patterns within existing datasets are often unclear or inadequately characterized [58]. Although fairness evaluation datasets aim to isolate specific biases, the nature and extent of these biases can remain ambiguous or obscured due to inconsistent definitions, subtle semantic structures, or limited interpretability methods [59]. Such ambiguity complicates researchers’ ability to precisely understand, compare, or effectively mitigate biases across different datasets. Therefore, a systematic survey of recent advances is needed to shed light on future research directions; however, no existing survey currently addresses this critical requirement.
@@
To this end, this paper aims to fill such a critical gap in the fairness literature by offering structured guidance on how to choose, examine, and refine datasets for evaluating large language models—datasets that function not only as diagnostic tools for uncovering bias but also as essential building blocks for developing mitigation approaches. To the best of our knowledge, this paper represents the first comprehensive work to specifically survey and analyze datasets designed for evaluating and mitigating fairness issues in Language Models. Specifically, we propose a unified taxonomy that organizes existing datasets and introduce a bias analysis methodology to support the systematic identification of dataset-level biases. Together, these contributions support informed dataset selection and the development of more effective fairness interventions. The main contributions of this survey paper are summarized as:
@@
    Systematic Organization of Fairness Benchmarks. A unified taxonomy is proposed to systematically organize existing fairness corpora, offering a structured framework for understanding how dataset design choices influence the trade-off between control and real-world relevance.
@@
    Unified Framework for Evaluating Dataset-Level Bias. A unified analysis pipeline is introduced to enable consistent, task-agnostic evaluation of dataset-level biases. This framework supports reproducible, comparable assessments across diverse fairness benchmarks by aligning bias types with general-purpose statistical estimators and applying standardized evaluation procedures tailored to different dataset structures.
@@
Connection to existing surveys. A number of surveys have examined fairness in language models from different angles. For example, Li et al. [60] provide a comprehensive review of fairness research in LLMs, offering separate taxonomies for medium-sized and large-sized models. Gallegos et al. [59] propose a taxonomy of bias evaluation and mitigation techniques, while Chu et al. [23] curate resources related to bias evaluation. Blodgett et al. [61] deliver a critical survey of bias in NLP more broadly. While these surveys offer valuable perspectives on fairness and bias in LMs, they largely treat datasets as supporting elements rather than subjects of focused analysis. In particular, none provide a consolidated and systematic examination of fairness datasets—their construction, characteristics, limitations, and appropriate applications.
This gap motivates our survey, which distinguishes itself by centering datasets as a primary object of analysis. We consolidate a wide range of fairness-related datasets scattered across the literature and organize them within a unified taxonomy. Our framework extends prior efforts by offering a multi-dimensional analysis that not only captures structural and linguistic properties but also incorporates a principled bias assessment methodology. This enables researchers to make more informed decisions when selecting datasets aligned with their model architectures, linguistic settings (e.g., monolingual or multilingual), and fairness priorities (e.g., focusing on specific protected attributes or evaluating multiple bias types).
Survey Structure. The remainder of this paper is organized as follows. Section 2 introduces the general experimental protocol that maps each bias type to a reproducible statistic, formalizes core bias concepts, and establishes a unified framework for measurement. Based on this foundation, Section 3 analyzes counterfactual benchmarks using task-specific evaluation procedures. Section 4 extends the analysis to prompt-based datasets applying the same methodological lens. Section 5 synthesizes insights from both structural families to offer practical guidance on selecting and combining benchmarks for comprehensive fairness evaluation. Finally, Section 6 concludes the survey and outlines directions for future work.
@@ 2 Taxonomy and Analysis Methodology
We first outline the dataset search and screening procedure that reduced 137 candidates to 16 representative benchmarks. Figure 1 then summarises our unified taxonomy: fairness datasets are grouped into two structural families—counterfactual and prompt-based—and situated within five intersecting dimensions (structure, source, linguistic coverage, bias typology, and accessibility). Building on this taxonomy, we introduce a bias analysis methodology that formalises four types of dataset-level bias—representativeness, annotation bias, stereotype leakage, and differential metric bias—through principled statistical estimators.
@@
2.1 Dataset Collection and Selection Criteria
2.1.1 Search Strategy
To identify relevant datasets and benchmarks, major academic databases (IEEE Xplore, ACM Digital Library, Scopus, Google Scholar) were searched using “fairness datasets” and ‘fairness benchmarks” as primary terms, supplemented by secondary terms such as “bias,” “discrimination,” and “language models.” This strategy accounted for the interdisciplinary nature of NLP fairness research, which intersects AI, ethics, linguistics, and social computing. The review prioritized publications from leading conferences and journals focused on datasets, benchmarks, NLP, and language models, including ACL [62, 63, 64, 65, 66, 67, 68, 69, 70, 57], EMNLP [71, 72, 56, 73, 74, 75], NAACL [76], and ACM FAccT [77]. Additionally, references from selected papers and existing surveys on fairness in language models [23, 78, 59, 29] were examined to ensure comprehensive coverage.
2.1.2 Inclusion and Exclusion Criteria
Inclusion Criteria. A dataset was admitted into the corpus if it primarily targets the detection of unjustified adverse effects on specific individuals or social groups, offers a transparent and replicable account of its construction, and demonstrates practical relevance through peer citations or real-world deployments. We further required comprehensive technical documentation—including data format, label taxonomy, and annotation procedures—to ensure reproducibility, as well as an explicit connection to current research questions in LM fairness. Only datasets satisfying all of these conditions advanced to the analysis stage.
From an initial search that identified 137 potential datasets, these criteria allowed for a focused selection, narrowing the list to 16 datasets that fully met the requirements for evaluating fairness in language models.
2.2 Datasets Taxonomy
Building on the dataset search and selection protocol described in Section 2.1, we now introduce a five-dimensional taxonomy that underpins the remainder of our analysis. The taxonomy organises fairness datasets by structure, source, linguistic coverage, bias typology, and accessibility—dimensions chosen to capture both how the data are constructed and how they should be applied. Mapping each benchmark to these interconnected axes provides researchers with a systematic way to locate resources that match their language-model setting and specific fairness objectives.
2.2.1 Structure
We categorize fairness datasets by their data structure into two major categories: (1) counterfactual datasets and (2) prompt-based datasets. This classification captures not only the structural differences but also also aligns closely with distinct evaluation methodologies used to measure fairness in LMs. Counterfactual datasets evaluate fairness by measuring how models respond differently when only identity-related attributes or stereotypes are varied in otherwise identical contexts, while prompt-based datasets evaluate fairness by analyzing biases manifested in model-generated content. This taxonomy provides researchers with a clear framework for navigating available fairness evaluation resources and select appropriate datasets based on their specific research objectives.
Counterfactual Datasets. Counterfactual datasets evaluate LM fairness by comparing responses to nearly identical sentences that differ only in identity-related attributes or stereotypes. By isolating the impact of these characteristics, researchers can accurately detect biases in scenarios where demographic features should ideally have no influence. Such evaluations are valuable for ensuring fairness in applications like coreference resolution, language generation, sentiment analysis, and classification tasks.
Based on the nature of their target evaluation tasks and linguistic manipulation strategies, counterfactual datasets can be subdivided into three categories: (1) Coreference Resolution Bias Datasets, (2) Stereotype Association Datasets, and (3) Sentiment and Classification Bias Datasets. Each subcategory addresses fairness from distinct perspectives, though they all share the fundamental approach of employing controlled pairs or minimal variations in their structure.
i) Coreference Resolution Bias Datasets. Coreference resolution bias datasets focus explicitly on biases in pronoun resolution tasks. They typically contain pairs of sentences that differ solely in pronouns referring to entities with varied demographic attributes (commonly gender). Examples of datasets in this category include WinoBias [62] and Winogender Schemas [79]. By analyzing differences in model predictions across these minimally contrasted sentence pairs, researchers can precisely quantify the biases related to gender associations in coreference tasks.
ii) Stereotype Association Datasets. Stereotype association datasets examine how stereotypes or biased associations toward various demographic groups affect language model outputs. These datasets comprise pairs or tuples of sentences where identity-related characteristics, such as gender, race, nationality, or socio-economic status are altered systematically to contrast stereotypical and anti-stereotypical associations. Examples include CrowS-Pairs [56] and French CrowS-Pairs [80]. By comparing model predictions on stereotypical versus anti-stereotypical sentences, researchers can directly identify biases embedded within model-generated probabilities or responses.
iii) Sentiment and Classification Bias Datasets. Sentiment and classification bias datasets employ controlled sentence templates or minimal pairs specifically tailored to evaluate biases in sentiment analysis and text classification tasks. By systematically varying only demographic descriptors within otherwise identical sentiment-bearing or classification-related contexts, datasets like the Equity Evaluation Corpus (EEC) [76] reveal how identity attributes unjustifiably influence model predictions. For instance, different sentiment scores assigned to “I am a Black woman and I love this song” versus “I am a White man and I love this song” would signal an underlying bias in a sentiment model.
Overlap Between Subcategories. While these subcategories distinctly delineate evaluation objectives, they inherently share methodological foundations and thus exhibit conceptual overlaps. The common characteristic among all three subcategories is their reliance on minimal linguistic pairs or controlled textual variations, in which only identity-related attributes are manipulated. For example, both Coreference Resolution Bias Datasets and Stereotype Association Datasets frequently involve sentences differing primarily by demographic markers, such as gender pronouns or occupational stereotypes. Consider the sentence pairs used in WinoBias [62] (a coreference-focused dataset): “The doctor asked the nurse to help her” versus “The doctor asked the nurse to help him.” Similar minimal-pair constructions appear in stereotype-focused datasets like CrowS-Pairs [56], e.g., comparing “The Black man was aggressive” versus “The White man was aggressive.” In both cases, the key technique, varying a single demographic indicator to measure model bias remains consistent, despite differences in the primary evaluation goal.
Similarly, overlaps exist between Stereotype Association Datasets and Sentiment and Classification Bias Datasets, as both categories manipulate identity attributes to examine whether model predictions unfairly shift due to stereotypes or biases. For instance, a sentiment-oriented dataset like the Equity Evaluation Corpus (EEC) [76] uses variations such as “I am a Black woman, and I love this song” versus “I am a White man, and I love this song.” Comparable minimal identity manipulations appear in stereotype datasets, such as sentence pairs that examine associations between occupations and demographic groups, e.g., “The software developer was a man” versus “The software developer was a woman.”
These overlaps highlight shared methodological principles across counterfactual datasets. However, despite structural similarities, each subcategory remains distinct due to its primary analytical objectives—whether evaluating bias in coreference resolution, stereotype associations, or sentiment and classification tasks. Understanding these nuances helps researchers select datasets aligned with their fairness evaluation goals while recognizing opportunities for cross-category methodological insights.
Prompt-based Datasets. Prompt-based datasets require the model to generate responses based on an initial prompt or question. They are subdivided into Sentence Completion and Question-Answering formats.
i) Sentence Completion Datasets. Generation-based datasets [81, 70, 77, 82, 83, 74] contain the starts of sentences, which can then be completed by the LM. Evaluations focus on detecting biases manifest in the unconstrained generation of text. Such datasets reflect realistic LM usage scenarios and provide a broader scope for capturing diverse forms of bias.
ii) Question-Answering Datasets. Question-answering datasets [57, 75, 84, 85] consist of structured queries designed to elicit model biases explicitly. These datasets often use ambiguous or underspecified contexts to test whether model answers align with stereotypes, effectively revealing biases in specific demographic assumptions.. This structured format facilitates clear interpretation and direct assessment of societal biases reflected in model responses.
2.2.2 Source
@@
The source of data plays a crucial role in determining the coverage, reliability, and representativeness of fairness evaluations in LMs. There are four primary sources of datasets used for this purpose: (1) template-based and external sources, (2) natural text sources, (3) crowdsourced sources, and (4) AI-generated sources.
@@
Template and external sources datasets [62, 71, 63, 64, 76, 75], combine predefined structures or patterns with an external source (e.g., US Department of Labor Labor Force Statistics from the Current Population Survey, 2017. https://www.bls.gov/cps/cpsaat11.htm) to test models for specific biases, especially in resolving ambiguities. These datasets are advantageous because they offer precise control over sentence variability and demographic terms, allowing for targeted fairness evaluations. However, their limitation lies in the fact that they are predefined, which reduces the linguistic diversity that could be encountered in natural language use.
@@
2.2.3 Linguistic Coverage
Linguistic coverage represents a critical dimension in fairness dataset taxonomy, determining the cross-cultural applicability and generalizability of fairness evaluations. This dimension primarily distinguishes between monolingual datasets (predominantly English) and those with broader multilingual scope.

### v2

@@ Abstract
Despite the growing reliance on fairness benchmarks to evaluate language models, the datasets that underpin these benchmarks remain critically underexamined. This survey addresses that overlooked foundation by offering a comprehensive analysis of the most widely used fairness datasets in language model research. To ground this analysis, we characterize each dataset across key dimensions, including provenance, demographic scope, annotation design, and intended use, revealing the assumptions and limitations baked into current evaluation practices. Building on this foundation, we propose a unified evaluation framework that surfaces consistent patterns of demographic disparities across benchmarks and scoring metrics. Applying this framework to sixteen popular datasets, we uncover overlooked biases that may distort conclusions about model fairness and offer guidance on selecting, combining, and interpreting these resources more effectively and responsibly. Our findings highlight an urgent need for new benchmarks that capture a broader range of social contexts and fairness notions. To support future research, we release all data, code, and results at https://github.com/vanbanTruong/Fairness-in-Large-Language-Models/tree/main/datasets, fostering transparency and reproducibility in the evaluation of language model fairness.
@@ 1 Introduction
Language Models (LMs), such as BERT [1], ELMo [2], RoBERTa [3], GPT-4 [4], LLaMA-2 [5], LLaMA-3 [6], and BLOOM [7], have demonstrated remarkable performance in various natural language processing tasks, including translation [8], text sentiment analysis [9], and text summarization [10]. This success has led to their widespread adoption across high-stakes domains such as hiring [11], loan approvals [12], legal sentencing [13], and medical diagnoses [14]. However, as their influence grows, a critical concern has emerged: most LMs were not designed with fairness considerations, which potentially leads to discriminatory outcomes in these downstream applications. For example, research has shown that these models can perpetuate stereotypes, linking leadership qualities with men while describing women as being “warm and friendly” [15]. Such biases in LMs raise significant ethical and societal concerns, underscoring the need for rigorous fairness research in language models.
@@
To address these fairness concerns, an expanding body of research focuses on understanding, assessing, and mitigating the biased behaviors exhibited by language models. Datasets are central to these efforts, serving not only as datasets for identifying fairness issues but also as the critical foundations for designing effective mitigation strategies [16]. For instance, WinoBias by Zhao et al. [17], which uses controlled sentence structures to isolate gender bias; CrowS-Pairs by Nangia et al. [18], which employs crowd-sourcing to capture naturally occurring stereotypes; and BBQ by Parrish et al. [19], which relies on expert annotation for high-quality bias measurements across nine demographic categories.
However, while dataset designs span tightly controlled templates to crowd-sourced naturalistic prompts, the landscape is fragmented and lacks a coherent organizational scheme. This heterogeneity affects study design and interpretation, creating two practical challenges for dataset selection [16]. First, goal alignment and comparability: differences in elicited outputs and evaluation metrics make it hard to choose benchmarks that match specific fairness goals and to compare results across studies. Second, opacity of bias signals: even when datasets target particular biases, their nature and extent often remain under-specified due to inconsistent definitions, subtle semantic structures, or limited interpretability [20, 21]. These gaps motivate a systematic survey that (i) formalizes core bias concepts with reproducible statistics, (ii) organizes datasets by output structure and complementary attributes, and (iii) distills selection and combination guidance for comprehensive fairness evaluation.
@@
In response to these challenges, this survey addresses the gap by offering structured guidance on how to choose, examine, and refine datasets for evaluating language models. We treat datasets not only as diagnostic instruments for revealing bias but also as foundational components for developing mitigation. To the best of our knowledge, this paper represents the first work to specifically survey and analyze datasets designed for evaluating and mitigating fairness issues in language models. Specifically, we propose a unified taxonomy that organizes existing datasets and introduce a bias analysis methodology to support the systematic identification of dataset-level biases. Together, these contributions support informed dataset selection and the development of more effective fairness interventions. The main contributions of this survey paper are summarized as:
@@
    Systematic Organization of Fairness Datasets. We propose a two-way taxonomy of open‑ended versus constrained-form evaluation datasets, which captures whether a model must freely generate text or choose among fixed outputs. This principled split clarifies how dataset format shapes fairness findings and guides dataset selection.
@@
    Unified Framework for Evaluating Dataset-Level Bias. A unified analysis pipeline is introduced to enable consistent, task-agnostic evaluation of dataset-level biases. This framework supports reproducible, comparable assessments across diverse fairness datasets by aligning bias types with general-purpose statistical estimators and applying standardized evaluation procedures tailored to different dataset structures.
@@
Connection to existing surveys. A number of surveys have examined fairness in language models from different angles. For example, Li et al. [22] provide a comprehensive review of fairness research in LLMs, offering separate taxonomies for medium-sized and large-sized models. Gallegos et al. [21] propose a taxonomy of bias evaluation and mitigation techniques, while Chu et al. [16] curate resources related to bias evaluation. Blodgett et al. [23] deliver a critical survey of bias in NLP more broadly. While these surveys offer valuable perspectives on fairness and bias in LMs, they largely treat datasets as supporting elements rather than subjects of focused analysis. In particular, none provides a consolidated and systematic examination of fairness datasets, including their construction, characteristics, limitations, and appropriate applications.
Survey Structure. The remainder of this paper is organized as follows. Section 2 introduces the general experimental protocol that maps each bias type to a reproducible statistic, formalizes core bias concepts, and establishes a unified framework for measurement. Based on this foundation, Section 3 examines constrained-form datasets, where models make fixed-choice decisions or rank alternatives. Section 4 analyzes open-ended evaluation datasets, where models generate text freely in response to prompts. Section 5 synthesizes insights from both structural families to offer practical guidance on selecting and combining datasets for comprehensive fairness evaluation. Finally, Section 7 concludes the survey and outlines directions for future work.
@@ 2 Taxonomy and Analysis Methodology
This section establishes the conceptual and empirical foundation for our survey. We first introduce a hierarchical taxonomy that situates fairness datasets within two structural families and four orthogonal attribute dimensions, providing a clear map for navigating the landscape of available resources. We then describe our dataset collection and screening pipeline, which yields the representative set of datasets analyzed in this work. Finally, we define three major forms of dataset-level bias and present task-agnostic estimators for each, forming a unified measurement framework that is applied consistently throughout Sections 3 and 4.
2.1 Datasets Taxonomy
Figure 1 presents our two-level taxonomy. At the first level, datasets are categorized into two structural families: constrained-form, where models choose or rank from fixed outputs, and open-ended, where models generate free-form text. This distinction captures the fundamental difference in model interaction and also organizes the remainder of this section: constrained-form datasets are detailed in Section 3, and open-ended datasets are detailed in Section 4. At the second level, each dataset is further described by four orthogonal attribute dimensions: sources, linguistic coverage, bias typology, and accessibility. This taxonomy provides researchers with a clear framework for navigating fairness evaluation resources, enabling them to select appropriate datasets based on their specific research objectives and the output structure they wish to evaluate.
@@ 2.1 Datasets Taxonomy
2.1.1 Structural Families
2.1.1.1  Constrained‑form Evaluation Datasets
Constrained-form evaluation datasets require the model to select or rank from a predefined set of outputs. These datasets are designed to assess fairness in specific decision-making or selection-based tasks, where the model is constrained by a limited output space, such as classification labels, multiple-choice answers, or ranking orders. The key advantage of this structure is the clear, quantifiable nature of the bias measures it supports, such as accuracy gaps, selection rates, or probability shifts across different demographic groups. We further organise this family into five task‑oriented sub‑categories.
i) Coreference and Pronoun Resolution Datasets. Here the model must decide or rank whether pronouns match their antecedents. Standard metrics include overall accuracy, F1, and the error‑rate gap between demographic conditions. WinoBias, WinoGender, and GAP exemplify this setting, each providing minimally contrasted sentence pairs that differ only in gendered or identity‑bearing pronouns.
ii) Sentence-Likelihood Counterfactuals Datasets. Datasets in this sub‑category present sentence pairs that differ solely by an identity term, allowing log‑likelihood or perplexity to serve as a fairness probe. The primary statistic is the change in perplexity ($\Delta\text{PPL}$) across protected groups. StereoSet, CrowS‑Pairs, RedditBias, and HolisticBias belong here, revealing subtle shifts in a model’s probabilistic preferences when demographic markers are swapped.
iii) Classification-based Bias Datasets. These datasets cast bias detection as a supervised prediction task, including sentiment analysis, natural-language inference, or toxicity classification, and then measure disparities such as group‑specific accuracy gaps or probability differentials. The Equity Evaluation Corpus and Bias NLI allow researchers to isolate whether models assign different class probabilities to semantically similar inputs that vary only in demographic references.
iv) Multiple-Choice Question Answering Bias Datasets. In multiple‑choice settings the model selects one option from a fixed set, where one candidate often encodes a stereotype or biased inference. Evaluation relies on bias scores or differential selection rates between protected attributes. Datasets such as BBQ and UnQover leverage underspecified or ambiguous questions to expose stereotypical reasoning in model choices.
v) Information Retrieval Bias Datasets. Datasets in this subcategory focus on evaluating whether ranking or similarity scores exhibit demographic parity. These datasets assess how well models balance fairness when ranking items, such as documents, products, or other entities, in response to a query. Common fairness metrics include the gap in Mean Reciprocal Rank (MRR) and normalized Discounted Cumulative Gain (nDCG), which quantify how ranking decisions differ across demographic groups. Grep‑BiasIR is a key example of this subcategory, providing datasets to audit bias in information retrieval and recommendation systems, where demographic factors may influence the ranking of results or selections.
2.1.1.2  Open-Ended Evaluation Datasets
In contrast, open-ended evaluation datasets require the model to generate text freely in response to prompts or instructions, without predefined output constraints. These datasets are particularly useful for analyzing emergent biases in unconstrained content generation, such as stereotypical associations, toxicity, or social alignment in the model’s outputs. The key challenge with open-ended datasets is that they assess the biases present in model-generated content, which can be influenced by subtle and complex patterns in the data.
Open-ended datasets do not rely on fixed labels or outputs but instead examine how the model responds to open-ended queries. Common tasks include text generation, completion, and summarization, where the model is expected to provide coherent and contextually relevant responses. Datasets such as BOLD, RealToxicityPrompts, HONEST, and TrustGPT are examples of open-ended evaluation datasets. These datasets allow for a broad examination of various biases in the model’s language generation, focusing on issues such as toxicity, sentiment, and fairness across different social or demographic groups.
2.1.2 Attribute Dimensions
Where the structural families defined above specify how a dataset constrains model outputs, a complementary set of attribute dimensions specifies what kinds of data underlie those structures and for whom the resulting fairness claims hold. We identify four orthogonal axes that collectively shape a dataset’s empirical validity and practical utility: Source, Linguistic Coverage, Bias Typology, and Accessibility. Source determines the degree of experimental control versus ecological realism; linguistic coverage delimits the cross-cultural scope of any fairness conclusion; bias typology specifies the protected attributes or construction artefacts being probed; and accessibility governs the ease with which the community can verify, replicate, and extend reported results. The remainder of this subsection elaborates each dimension in turn.
2.1.2.1  Source
@@
Having delineated the two structural families of fairness datasets, we now turn to the source of the datasets themselves. The source of data plays a crucial role in determining the coverage, reliability, and representativeness of fairness evaluations in LMs. We distinguish four primary sources of datasets used for this purpose: (1) template-based and external sources, (2) natural text sources, (3) crowdsourced sources, and (4) AI-generated sources.
@@
Template and external sources datasets [24, 25, 26, 27, 28, 29], combine predefined structures or patterns with an external source (e.g., US Department of Labor  [^1] Labor Force Statistics from the Current Population Survey, 2017. https://www.bls.gov/cps/cpsaat11.htm) to test models for specific biases, especially in resolving ambiguities. These datasets are advantageous because they offer precise control over sentence variability and demographic terms, allowing for targeted fairness evaluations. However, their limitation lies in the fact that they are predefined, which reduces the linguistic diversity that could be encountered in natural language use.
@@
2.1.2.2  Linguistic Coverage
Linguistic coverage represents a critical dimension in the fairness dataset taxonomy, determining the cross-cultural applicability and generalizability of fairness evaluations. This dimension primarily distinguishes between monolingual datasets (predominantly English) and those with broader multilingual scope.